1. The decomposition of N2O5 to NO2 and O2 is first order, with a rate constant of 4.8 x 10 -4 s -1 .

a) if the initial conc is 1.65 x 10 -2 M , what is the conc after 825 s?

b) How long would it take for the conc of N2O5 to decrease to 1 x 10 -2 M?
1
Expert's answer
2017-10-11T11:35:07-0400
2N2O5 = 2NO2 + O2
k = 4.8*10^-4 s^-1 (1 order reaction)
C = C0*e^-kt
lnC = lnC0-kt
C0 = 1.65*10^-2 M


a)t=825s
C = 1.65*10^-2*e^(-4.8*10^-4*825)= 1.65*10^-2*0.673 = 1.11*10^-2

b)
lnC = lnC0-kt
ln(1*10^-2)=ln(1.65*10^-2)-4.8*10^-4*t
t= -ln((1*10^-2)/(1.65*10^-2))/4.8*10^-4 = (0.5/4.8)*10^4 = 1041s